Below are 4 perfect competition firms who want to max their profits. Each row represents a separate firm. You have been hired as a consultant to assist them in doing so. Please advise them as to what decisions they should make to accomplish this by using the rules of marginal analysis for PC firms. See above for the due date for this assignment.

1. S hould the firm increase or decrease output, or leave output unchanged? How do you know? 

2. Can the firm make a profit, earn a normal profit, minimize a loss, or should the firm shutdown? How do you know?

3. What happens to price if output changes?

 Be sure to state reasons for your recommendations.  

PC firms

 

Price     Q            TR          TC          P/L        TVC       ATC       AVC      MC

4             100        400        350        +50        300        3.5         3             5

10          20          200        500        -300      300        25          15          10          

50          100        5000      5100      -100      3000      51          30          90          

25          100        2500      2500      0             2000      25          20          25


1
Expert's answer
2021-02-18T07:13:54-0500

1. In a perfect competitive markets profits will be highest and losses small at the quantity where total revenues exceed total cost by a great amount.The p/l for the firms are 100,-300,-100 and 0 respectively. The firm with total revenue of 400 should leave the output unchanged because it's already making profit ,because total cost exceeds the total revenue . All the other firms with total revenue of 200,5000,2500 should increase their output so as to maximize their profit. This is bacause an increase in output will lead to increased total revenue.


2. "Profit =TR-TC"

"=400-350=50"

This firm is making profit.

"=200-500=-300"

This firm is making losses. So the firm should increase quantity of the output.

"=5000-5100=-100"

The firm is making losses. So the firm should increase quantity of the output.

"=2500-2500=0"

No profits. So the firm should increase quantity of the output.

3. In a competitive market , when demand increases the price for the goods increase too. Thus an increase of output will lead to an increase in prices of the commodities.